OpenStudy (anonymous):

Simplify 216^1/3. Answer 4 6 9 12

OpenStudy (anonymous):

216^1/3 implies the cube root of 216, so what number when multiplied by itself 3 times you get 216?

OpenStudy (anonymous):

6 :) @LAGIRL69

OpenStudy (zepp):

\[\LARGE x^{\frac{1}{2}} =\sqrt{x}\\\LARGE x^{\frac{1}{3}}=\sqrt[3]x\\\LARGE x^{\frac{1}{4}}=\sqrt[4]{x}\] :)

OpenStudy (anonymous):

Thanks Thank

OpenStudy (zepp):

@Tank http://openstudy.com/code-of-conduct Try to not give the answer right away, try to help her understand the concept instead :)

OpenStudy (anonymous):

im not a good teach plus i suck at math lol

OpenStudy (anonymous):

simplify (m3/3n)/(-m4 /9n2) like this what is this crap

OpenStudy (anonymous):

@Tank can you relate your expression to this? \[(a/b) / (-a/-b) = (a/b) \div (-a/-b) = a/b * -b/-a\]

OpenStudy (anonymous):

oh dear

OpenStudy (zepp):

It doesn't have to be negative on the downstairs

OpenStudy (zepp):

\[\large \frac{a}{b}\div\frac{c}{d}=\frac{a}{b}*\frac{d}{c}\]

OpenStudy (anonymous):

well am i just adding like terms

OpenStudy (anonymous):

cause i keep geting m/3m2

OpenStudy (zepp):

\[\large \frac{m^3}{3n}\div\frac{-m^4}{9n^2}=\frac{m^3}{3n}*\frac{9n^2}{-m^4} \]

OpenStudy (zepp):